OpenStudy (slickster256):

Which of the following sets shows all the numbers from the set {0.5, 1, 2.5, 3, 3.5} that make the inequality 4a + 2 ≥ 12 true? {2, 2.5, 3} {2.5, 3, 3.5} {0.5, 1, 2.5} {1, 2.5, 3}

OpenStudy (slickster256):

plz help i dont under stand

OpenStudy (misty1212):

lets try the first one is \[4\times 2+2\geq 12\]?

OpenStudy (slickster256):

10

OpenStudy (slickster256):

so true

OpenStudy (misty1212):

really? ten is greater than or equal to twelve?

OpenStudy (slickster256):

no its not

OpenStudy (misty1212):

ok so A is the the answer

OpenStudy (slickster256):

ok i need a lil more help

OpenStudy (misty1212):

and so neither is C or D, because both of those start with numbers even less than 2

OpenStudy (misty1212):

i meant "A is NOT the answer"

OpenStudy (slickster256):

ok do it would be b

OpenStudy (misty1212):

yes B works you can check

OpenStudy (misty1212):

\[4\times 2.5+2=10+2=12\]and \[12\geq 12\] is true

OpenStudy (slickster256):

Which of the values in the set {1, 2, 3, 4} is a solution to the equation 3x + 3 = 15? 1 2 3 4

OpenStudy (slickster256):

@misty1212

OpenStudy (misty1212):

\[3x+3=15\] subtract 3 get \[3x=12\] then divide by 4

OpenStudy (slickster256):

ok

OpenStudy (slickster256):

i think i got tht

OpenStudy (slickster256):

so its D

OpenStudy (misty1212):

it is 4

OpenStudy (slickster256):

ok thx have a nice day misty

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]